OpenStudy (anonymous):

Let f(x)=x^3-7x^2+25x-39 and let g be the inverse function of f. What is the value of g'(0)?

OpenStudy (anonymous):

Since f(3)=0, we know that g(0)=3. The slope of the inverse function g is given by the formula \[g'(x)=1/f'(g(x))\]. Thus, \[g'(0)=1/f'(g(0))=1/f'(3)\]. Since \[f'(x)=3x^2-14x+25\], we know that \[f'(3)=3*3^2-14*3+25=10\]. Thus, \[g'(0)=1/10.\]

OpenStudy (anonymous):

Thanks can you help with another?

OpenStudy (anonymous):

Sure, if I'm able to help, I will. What's the question?